CourtneyWatson:

Myra owns a car service that charges a $5 flat rate and an additional $0.50 per mile. 1. y−intercept = (0, 5), point = (x, y) 2. m = 0.5 3. 0.5 = y − 5 x − 0 →0.5 = y − 5 x 4. Solve for y : 0.5 = y − 5 x →0.5x = y − 5

jakfishman:

mxddi3 could prob help

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jakfishman mxddi3 could prob help \(\color{#0cbb34}{\text{End of Quote}}\) Ok

CourtneyWatson:

could you help me please @mxddi3

supie:

@Mercury maybe he can help....

mxddi3:

uhm ok so i think we just have to get an equation in y=mx+b form ? am i correct?

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 uhm ok so i think we just have to get an equation in y=mx+b form ? am i correct? \(\color{#0cbb34}{\text{End of Quote}}\) you have to solve for Y

supie:

Thats what she is saying, if y=mx+b then we are solving to get the value of y but first you have to plug in the numbers i think.

mxddi3:

but like, if you put it into slope-intercept form, you need to add 5 to both sides and from there, you will have your equation as y=mx+ b do you have a point we can plug in for this to solve this? lol supie you can help if u want as well.

supie:

idk it tho XD

mxddi3:

can you tell me what numbers i'm supposed to plug into the equation then? I am sorry, i'm slightly confused

CourtneyWatson:

um im confused too I looked it up on Brainly and this is what it said


>https://brainly.com/question/14371996

mxddi3:

ik how to get you the equation.. you have 0.5x=y-5 so you add 5 to both sides and -5+5 cancels out leaving you with y=0.5x+5 what else do you need?

CourtneyWatson:

so the answer is y=0.5x+5

CourtneyWatson:

i think

mxddi3:

i mean, i thought that as well....

CourtneyWatson:

Ok I got it right thank you so much for helping @mxddi3 and @supie

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CourtneyWatson so the answer is y=0.5x+5 \(\color{#0cbb34}{\text{End of Quote}}\) I believe that is the answer because the constant is 5 which is the y intercept/flat rate, and the .50 per mile is the changing variable

ramen:

Also, .5x=y-5, you would just add 5 to the left side making it y=.5x+5. So yes that was the right answer.

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen Also, .5x=y-5, you would just add 5 to the left side making it y=.5x+5. So yes that was the right answer. \(\color{#0cbb34}{\text{End of Quote}}\) Ok thank you

jimthompson5910:

y = mx+b is slope intercept form x = number of miles y = total cost m = 0.5 = cost per mile = slope b = 5 = flat rate = y intercept or initial value y = mx+b turns into y = 0.5x+5

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 y = mx+b is slope intercept form x = number of miles y = total cost m = 0.5 = cost per mile = slope b = 5 = flat rate = y intercept or initial value y = mx+b turns into y = 0.5x+5 \(\color{#0cbb34}{\text{End of Quote}}\) Thank you

jimthompson5910:

no problem